# The Rlign Algorithm for Enhanced Electrocardiogram Analysis through R-Peak Alignment for Explainable Classification and Clustering

### v1

@@ 2.1 R-peak alignment pipeline
The alignment comprises two main steps: initially detecting R-peaks and subsequently resampling to the form of a predetermined uniform ‘template’. The entire process is schematically visualized in Fig. 1. Our framework aligns with Scikit-Learn API conventions [11] and features an efficient multiprocessing pipeline to rapid processing of multiple ECGs.
@@ 2.2.4 Unsupervised Machine Learning
The analysis methodically evaluated the clustering performance using both the raw time series and various forms of realigned data. Throughout these tests, we consistently used a predefined cluster count of five, in alignment with the PTB-XL superclasses definitions. The performance was measured using the V-measure score, which is the harmonic mean between homogeneity and completeness, comparing the clustering results to the expert-labeled superclasses.
@@ 3.2 Classification performance and calibration
Our analyses reveal that on unaligned data, only the CNN performed with reasonable performance, whereas the LR and SVM hardly surpassed the level of chance in both binary and multi-class classifications. Alignment of ECG signals with median beat calculation or resampling to aligned full-length ECGs enabled similar performances of LR and SVM as compared to the CNN, with the SVM even outperforming the CNN in most tested scenarios. Within our experimental setup, the highest predictive performance was achieved with a SVM on Rlign hrc-resampled median beats.
@@
The analysis of model calibration reveals a distinct advantage for both shallow learning algorithms over the CNN when limited training data is available in binary classifications. The ECE, which quantifies the difference between predicted probabilities and actual outcomes (lower is better), for the CNN model exceeds 0.2 with half a fold of training data, whereas LR and SVM maintain an ECE below 0.05. This lower ECE indicates more reliable and accurate probability predictions from LR and SVM in the presence of limited training data. With three or more training folds, the performance of all algorithms levels out. In multiclass scenarios, however, the SVM notably outperforms the others, with LR closely matching CNN’s performance. In supplement 4.2, we provide a detailed analysis of the calibration for the Norm vs. STTC classification using a single fold of training data to demonstrate the differences in model probabilities.
@@ 3.3 Effects of R-peak alignment on unsupervised machine learning
Further analysis focused on differentiating the principal components associated with distinct compartments of the ECG, specifically the QRS complex and the T-wave, illustrating how R-peak alignment facilitates a more nuanced interpretation of ECG data. For Uni-G median beats (Fig. 4e), as expected, the normal ECGs and ST-T changes primarily differ in the T-wave component. However, the QRS component, crucial for identifying MI, fails to show differentiation, resulting in an overlap across all superclasses in this dimension. With Rlign processed signals (Fig. 4f-g), not only do normal ECGs and ST-T changes diverge in the T-wave component, but MI also shows a distinct pattern. Unlike ST-T changes, MI diverges less in the T-wave component but significantly in the QRS component, aligning with clinical expectations. This analysis highlights the advantage of R-peak alignment for enhancing the explainability of unsupervised learning in ECG analysis. By aligning ECG signals, Rlign allows for a clear differentiation of principal components associated with different ECG compartments and their distribution across diagnostic superclasses.
@@ 4 Discussion
All simulations are based on the publicly accessible PTB-XL data set. The Rlign source code is available under the terms of the MIT license at GitHub (https://github.com/imi-ms/rlign).
@@ 4.1 Heart-rate-corrected (hrc) Alignment
The various phases of the electrocardiogram (ECG) exhibit distinct dependencies on the underlying heart rate. In Fig. 2 of the main text, we illustrate the effectiveness of our method across bradycardia, tachycardia, and normal heart rate conditions. Our approach involves constructing a heart-rate-corrected (hrc) model that aims to accurately reconstruct the template frequency. Although existing literature explores this topic, no precise formula has yet been identified that describes the T-wave and P-wave as functions of heart rate. In further analysis, we utilized the PTB-XL dataset, categorizing ECGs into groups based on heart rates ranging in increments of 5 bpm steps from 50 to 105. Using Neurokit2, we computed both P-Onset and T-Offset, then normalized these measurements across the RR interval length to assess the proportion of phases within the ECG cycle. The findings of this analysis are presented in Fig. S1. The hrc-resampling formula is a linear approximation of this. Thus, the expected event points P-Onset and T-Offset are adjusted based on the initial measured heart rate $hr_{\text{measurement}}$, the template heart rate $hr_{\text{template}}$, and the RR interval. In detail, we resample three parts of every ECG cycle with different scales. First the interval between R-peak and T-offset
@@ 4.2 Calibration
In the medical field, attributes such as good calibration are gaining importance alongside the pure performance of machine learning models. To evaluate a model’s calibration, we divide the possible confidence interval $[0,1]$ into $M$ bins of size $\frac{1}{M}$. Let $B_{m}$ represent the set of sample indices where the prediction confidence falls within the interval $I_{m}=(\frac{m-1}{M},\frac{m}{M}$]. The accuracy of $B_{m}$ is defined by

### v2

# suppSupplementary References 1 2 3 * 1,2,* 1,* 1 1 1 3 1,2,3,* 1,*

@@ 2.1 R-peak alignment pipeline
The alignment comprises two main steps: initially detecting R-peaks and subsequently resampling to the form of a predetermined uniform ‘template’. The entire process is schematically visualized in Fig. 1. Our framework aligns with Scikit-Learn API conventions [16] and features an efficient multiprocessing pipeline for rapid processing of multiple ECGs.
@@ 2.2.4 Unsupervised Machine Learning
The analysis methodically evaluated the clustering performance using both the raw time series and various forms of realigned data. Throughout these tests, we consistently used a predefined cluster count of five, in alignment with the PTB-XL superclass definitions. The performance was measured using the V-measure score, which is the harmonic mean between homogeneity and completeness, comparing the clustering results to the expert-labeled superclasses.
@@ 3.2 Classification performance and calibration
Our analyses reveal that on unaligned data, only the CNN performed with reasonable performance, whereas the LR and SVM hardly surpassed the level of chance in both binary and multi-class classifications. Alignment of ECG signals with median beat calculation or resampling to aligned full-length ECGs enabled similar performances of LR and SVM as compared to the CNN, with the SVM even outperforming the CNN in most tested scenarios. Within our experimental setup, the highest predictive performance was achieved with an SVM on Rlign hrc-resampled median beats.
@@
The analysis of model calibration reveals a distinct advantage for both shallow learning algorithms over the CNN when limited training data is available in binary classifications. The ECE, which quantifies the difference between predicted probabilities and actual outcomes (lower is better), for the CNN model exceeds 0.2 with half a fold of training data, whereas LR and SVM maintain an ECE below 0.05. This lower ECE indicates more reliable and accurate probability predictions from LR and SVM in the presence of limited training data. With three or more training folds, the performances of all algorithms level out. In multiclass scenarios, however, the SVM notably outperforms the others, with LR closely matching CNN’s performance. In supplement 4.2, we provide a detailed analysis of the calibration for the Norm vs. STTC classification using a single fold of training data to demonstrate the differences in model probabilities.
@@ 3.3 Effects of R-peak alignment on unsupervised machine learning
Further analysis focused on differentiating the principal components associated with distinct compartments of the ECG, specifically the QRS complex and the T-wave, illustrating how R-peak alignment facilitates a more nuanced interpretation of ECG data. For Uni-G median beats (Fig. 4e), as expected, the normal ECGs and ST-T changes primarily differ in the T-wave component. However, the QRS component, crucial for identifying MI, fails to show differentiation, resulting in an overlap across all superclasses in this dimension. With Rlign processed signals (Fig. 4f-g), not only do normal ECGs and ST-T changes diverge in the T-wave component, but MI also shows a distinct pattern. Unlike ST-T changes, MI diverges less in the T-wave component but clearly in the QRS component, aligning with clinical expectations. This analysis highlights the advantage of R-peak alignment for enhancing the explainability of unsupervised learning in ECG analysis. By aligning ECG signals, Rlign allows for a clear differentiation of principal components associated with different ECG compartments and their distribution across diagnostic superclasses.
@@ 4 Discussion
All simulations are based on the publicly accessible PTB-XL data set [26]. The Rlign source code is available under the terms of the MIT license at GitHub (https://github.com/imi-ms/rlign).
@@ 4.1 Heart-rate-corrected (hrc) Alignment
[…]
@@ 4.2 Calibration
[…]